OpenStudy (anonymous):

if you have a height of 12cm and a altitude of 7cm and a base thats 8cm. what is the area of the trapezoid??

OpenStudy (alexwee123):

what how can you have 2 heights? altitude=height .....

OpenStudy (anonymous):

i messed up on the drawing a bit

OpenStudy (alexwee123):

lol it kind of looks like a rhombus o.0 well if you insist it's a trapezoid then use this formula (Base length 1 +Base length 2)/2 *Height=Area so Area= (12+8)/2*7 Area=70cm squared

OpenStudy (anonymous):

yeah its a rhombus lol.

OpenStudy (alexwee123):

wait so it's not a trapezoid? i was only jk o.0

OpenStudy (anonymous):

well it said you had to find the area of the trapezoid lol && haha okayy! but thanks i got it right

OpenStudy (alexwee123):

k ^.^

OpenStudy (anonymous):

yupp!!!